OpenStudy (anonymous):

determine whether the function is odd, even or neither? k(x)=3-7(sqrt1+x^2)

OpenStudy (lukecrayonz):

\[3-\sqrt{1+x^2}\]

OpenStudy (anonymous):

\[3-7\sqrt{1+x^2?}\]

OpenStudy (anonymous):

???????? @lukecrayonz

OpenStudy (lukecrayonz):

Oh right sorry 1 sec

OpenStudy (lukecrayonz):

Ok so input -x for everytime you see x, and tell me what you get.

OpenStudy (lukecrayonz):

If f(–x) = f(x), all of the signs are the same, then the function is even. If you end up with the exact opposite of what you started with (that is, if f(–x) = –f(x), so all of the "plus" signs become "minus" signs, and vice versa), then the function is odd

OpenStudy (lukecrayonz):

And if you get neither answer, then it's neither! :P

OpenStudy (anonymous):

Ok h(x)=\[3-7\sqrt{1+(-x)^2}\]

OpenStudy (lukecrayonz):

So it's even.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

there is another one please

OpenStudy (lukecrayonz):

Sure :-)

OpenStudy (anonymous):

h(x)=x(4-x^2)

OpenStudy (lukecrayonz):

Odd :-)

OpenStudy (anonymous):

h(-x)=-x[4-(-x)^2]

OpenStudy (anonymous):

so....

OpenStudy (anonymous):

what next ??

OpenStudy (anonymous):

??????????????????

OpenStudy (anonymous):

@lukecrayonz

OpenStudy (lukecrayonz):

if f(–x) = –f(x), so all of the "plus" signs become "minus" signs, and vice versa), then the function is odd